OpenStudy (anonymous):

Which of the following equations is written in slope-intercept form? 2x + 4y = 8 4y = -2x + 8 y = -1/2x + 2 x = -2y + 4

OpenStudy (hoblos):

the slope-intercept form has a general formula : y= ax + b from the options only (y = -1/2x + 2) respect the general formula.. so the answer is y = -1/2x + 2